OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -15, -6, 3, 12,

OpenStudy (mandre):

The general equation for the n-th term is a+(n-1)d. Do you know what each letter stands for?

OpenStudy (anonymous):

no, i never paid attention in class and thats why im redoing it online :/

OpenStudy (mandre):

a is the first term and d is the difference. You should be able to work it out now.

OpenStudy (anonymous):

but what is the n-1?

OpenStudy (mandre):

You're supposed to provide the n-th term right?

OpenStudy (mandre):

Just simplify then you have your n-th term.

OpenStudy (anonymous):

so it would be -15+(n-1)-9?

OpenStudy (mandre):

Simplify more.

OpenStudy (anonymous):

an = -15 + 9(n - 1)

OpenStudy (mandre):

Almost done.

OpenStudy (anonymous):

thats one of the choices

OpenStudy (anonymous):

an = -15 + 9(n + 1) an = -15 x 9(n - 1) an = -15 + 9(n + 2) an = -15 + 9(n - 1)